You will need to generate a set of graphs to answer this question.

Concentration vs time data for the following decomposition reaction was collected at 300 K. The data is shown in the table below.

What is the rate constant for the reaction at this temperature (in units of min-1)?

H2O2(aq) → H2O(l) + 1/2 O2(g)

time (min) [H2O2](mol/L)
0 0.00191
500 0.00082
1000 0.00035
1500 0.00015
2000 0.00006
Answer #1

We know that for first order expression;

K * t = 2.303 * log (A0/At)

Where K is rate constant, A0 is initial concentration and At is concentration at time t.

Now,

K * 500 = 2.303 * log (0.00191/0.00082)

K = 0.00169 min^-1 OR 0.0017 min^-1 .... Answer